### v1

# Stateful Large Language Model Serving with Pensieve

Abstract.
Large Language Models (LLMs) have recently experienced great success, as evident in the widespread popularity of ChatGPT. Existing LLM serving systems are stateless across requests. Consequently, when LLMs are used in the common setting of multi-turn conversations, a growing log of the conversation history must be processed alongside any request by the serving system at each turn, resulting in repeated history processing.
In this paper, we design $Pensieve$, a system optimized for multi-turn conversation LLM serving. $Pensieve$ maintains the conversation state across requests by caching previously processed history to avoid duplicate processing. $Pensieve$’s multi-tier caching strategy can utilize both GPU and CPU memory to efficiently store and retrieve cached data. $Pensieve$ also generalizes the recent PagedAttention kernel to support attention between multiple input tokens with a GPU cache spread over non-contiguous memory. Our evaluation shows that Pensieve is able to achieve 1.51-1.95x throughput compared to vLLM and reduce latency by 60-75%.
1. Introduction
The world has recently witnessed the fast expansion of Large Language Models (LLMs). The most popular use of LLMs are for chatbots, with applications like ChatGPT demonstrating an astounding capability of chatting and answering questions like real humans. Other LLM-backed applications are also being applied to automate a variety of tasks, such as writing ads, responding to emails, doing literature review etc. As LLM continues its explosive growth, it is imperative to develop fast and efficient LLM serving systems.
An LLM is an autoregressive DNN model based on Transformer (transformer, ) architecture. The model iteratively predicts the next output token based on the current context which includes the sequence of input prompt tokens followed by output tokens generated in the previous iterations. LLMs require very expensive computation due to two factors. One, LLMs have huge parameter sizes (10s or 100s of billions) and have a trend of growing even larger. Two, LLMs need to support a large context size (2K to 32K tokens) to be useful. There has been much related work which improve the performance of LLM inference/serving from various angles, including better batching (orca, ), operation fusion (flash, ), better GPU memory utilization (vllm, ), faster output generation (speculative:leviathan, ; speculative:chen, ), low rank adaptation (lora, ) and quantization (qlora, ) (see §8). In contrast to these work, in this paper, we take a step back and examine inefficiencies that arise in the context of a specific but very popular LLM use case today, aka as a multi-turn conversational chatbot.
In the conversational setup, the user and the chatbot are engaged in a dialogue that may last many rounds. In order for the chatbot not to “lose memory” of what has been said so far when responding, the cumulative history of the dialogue must be part of the context for LLM’s autoregressive generation. As existing LLM serving systems are stateless across requests, clients send a growing log of conversation history alongside each new request as the input prompt to be processed from scratch. This causes much duplicate processing for multi-turn conversations.
How to avoid duplicate processing of the chat history? To do so, the serving system can save any previously processed context data. When new requests from the same conversation arrive, the saved context data can be re-used and subsequently augmented. This can be done with best effort. Essentially, the serving system is allowed to keep some cached state containing processed context across requests. Doing so enables the serving system to exploit the opportunity that when users are actively chatting with an AI chatbot, follow-up requests usually arrive within a reasonably short time period to leverage cached state.
Caching state across requests is straightforward at the high-level, but several challenges remain to make it really work. First, where to save the data? Keeping it in the GPU is the fastest, but is very constrained by the relatively small GPU memory size. Putting it on disk would incur much longer load latency, hurting the user experience. A two-tier caching solution spanning both GPU and CPU memory is promising, but care must be taken to cope with each tier’s capacity limit and to swap in/out the saved context data from/to the GPU without damaging performance. Second, how to re-use saved context data efficiently when processing a new request? When restoring a conversation’s context data from CPU to GPU, how to merge it with the conversation’s existing context that is already cached in the GPU? Furthermore, what has been saved might not be complete due to the cache limit. How to handle dropped cache?
In this paper, we design a stateful LLM serving system, called Pensieve, to address the aforementioned challenges. Pensieve saves a conversation’s processed context in a two-tier GPU-CPU cache. It evicts cached data to the next tier (or discards it) according to how long the conversation has been inactive and how large the conversation’s saved context has accumulated. The eviction is done at the granularity of a suffix of the conversation instead of the whole. Therefore, a conversation’s saved context might span both tiers of the cache and also be partially trimmed. Pensieve uses ahead-of-time swapping and pipelined transfer to overlap computation with the data movement between cache tiers. Dropped contexts are handled via recomputation. Evicting and restoring cause a conversation’s cached context to occupy non-contiguous GPU memory. We develop a new GPU kernel to compute attention (transformer, ) between multiple input tokens and cached context residing in non-contiguous memory, which lacks in existing LLM serving systems. Our kernel is a generalized version of the PagedAttention kernel in vLLM (vllm, ).
We have built the Pensieve and compared its performance against vLLM (vllm, ), a state-of-the-art serving system that does not cache state across requests. Experiments show that Pensieve is able to improve serving throughput by 1.51-1.95x and reduce normalized latency by 60-75%.
In summary, this paper makes the following contributions:
  • We identify a major inefficiency of existing LLM serving systems when used for multi-turn conversations: a conversation’s history context is recomputed with each successive new request in the same conversation.
  • We develop Pensieve, a stateful LLM serving system that saves the conversation context in a multi-tier GPU-CPU cache and reuses it across requests to minimize redundant computation. Our system design can efficiently move data between cache tiers and handle partially dropped context via recomputation.
  • We build a new attention GPU kernel to compute attention between a new request’s multiple input tokens and the saved context scattered in non-contiguous GPU memory. Existing kernels either require contiguous GPU context cache or restrict to a single input token.
  • We evaluate Pensieve using real world conversation datasets to demonstrate its effectiveness compared to the state-of-the-art stateless serving system.
2. Background
Figure 1. Inference of Transformer-based Large Language Models
In this section, we provide a brief background on LLM, its core attention operation, and how existing systems serve LLMs.
2.1. LLM and the Attention Mechanism
Popular large language models, e.g. GPT-3 (gpt3, ), OPT (opt, ), Llama (llama, ), are all based on the Transformer architecture (transformer, ). A model consists of many transformer layers, each of which is composed of an attention module, seen as the dashed box Figure 1, and a 2-layer feed forward network. The model takes as input a sequence of token ids representing the natural language sentence and feeds them through an embedding layer to obtain a continuous representation (aka embedding) for each token before feeding them through transformer layers. For simplicity, we refer to token embeddings as “tokens”, and refer to token ids as “raw tokens”. LLM is autoregressive in the sense that it iteratively predicts the next output token based on the current context which includes the input prompt tokens followed by output tokens generated in the previous iterations.
The success of Transformer originates from the capability of its attention module. For each layer, the attention module first performs linear transformations on its input token embedding ($X$) to produce three new embeddings Query ($Q$), Key ($K$), and Value ($V$):
$Q$ $=XW_{query}$
(1) $K$ $=XW_{key}$
$V$ $=XW_{value}$
where $W_{query}$, $W_{key}$ and $W_{value}$ are trainable weights. We refer to the $K$ and $V$ embeddings as KV-tokens. The module then computes all-to-all attention scores using the dot product between each pair of tokens’ query and key:
(2) $A=\frac{QK^{T}}{scale}$
where scale is a normalization factor. This attention score is then normalized using softmax and used for weighted aggregation of the token embeddings in $V$:
(3) $O=softmax(A)V$
The above equations show the process of a so-called single-head attention. In practice, multi-head attention is used, where $Q$, $K$, $V$ produced by Equation (1) are divided into groups, called attention heads. Each attention head independently performs attention as Equation (2) and (3).
2.2. How LLM is Served
The prefill vs. generation phase
To perform inference using an LLM, one needs to keep a KV cache in GPU to avoid recomputation during the autoregressive output generation. Figure 1 shows the typical LLM inference process adopted by systems like FasterTransformer (fastertransformer, ), ORCA (orca, ), vLLM (vllm, ). It is divided into two phases: 1) The prefill phase processes all the input prompt tokens all together to generate the $K$ and $V$ token embeddings (aka KV-tokens) for each Transformer layer and initialize the KV cache with the resulting KV-tokens. The embedding of the last token of the prompt from the last layer’s $O$ tensor is used to generate the first output token; 2) The iterative generation phase takes in the token generated by last step as a single new input token. For each layer, it computes the $Q$,$K$,$V$ embedding for the new token, updates the KV cache, and performs attention using the new token’s $Q$ embedding with the entire context’s KV-tokens.
Iteration-level batching
For LLMs have variable sized input and output, batching granularity has a huge impact on system throughput and serving latency. If scheduling is performed at the request granularity, executing a batch of requests with different input prompt lengths require padding tensors to the maximum length and waiting for the request with the longest output to finish. Iteration-level batching strategy, originally proposed by BatchMaker (batchmaker, ) for non-transformer-based sequence-to-sequence models, performs batching at token granularity. Whenever a request finishes an iterative generation step, the scheduler checks whether it has reached the end of sequence and can leave the batch, making room for a request to start its computation immediately. As we advance into Transformer era, ORCA (orca, ) extends this approach to support the LLM workload.
Memory management.
For each request, the model performs iterative generation until either the special end-of-sentence token (EOS) is emitted or the preconfigured maximum decoding length is reached. Systems like FasterTransformer (fastertransformer, ) and ORCA (orca, ) reserve slots in KV cache for each request based on the maximum decoding size. A more recent system, vLLM (vllm, ), can dynamically grow the allocated cache slots for each request and allow these slots to reside in non-contiguous GPU memory. vLLM develops PagedAttention to handle the generation phase with non-contiguous KV cache. Existing serving systems are stateless across requests. In other words, they de-allocate all the cache slots used by a request as soon as it finishes.
3. Motivation and Challenges
3.1. Motivation
Figure 2. Existing serving systems send a cumulative history with each request in a multi-turn conversation.
Existing techniques for serving Large Language Model mostly focus on improving the inference time or memory efficiency of a single request. We take a step back and examine inefficiencies that arise in the context of a specific but very popular LLM use case today, aka as a multi-turn conversational chatbot.
In a multi-turn conversation, user has multiple rounds of conversations with the chatbot and LLM needs to be aware of the conversation history in order to generate appropriate response. This is done by prefixing each new request with the entire cumulative conversation history, due to the stateless nature of existing serving systems, as shown in Figure 2. As the interaction between user and chatbot goes on, conversation history grows, making the cost of the prefill phase soon overshadow that of iterative generation phase. Unfortunately, much of the history processing is redundant.
Figure 3. Execution time for a batch of 32 requests performing prompt (32 tokens) prefill and generations for 200 steps.
The goal of this project is to minimize redundant computation of the conversation history. This can be done by caching any previously processed embeddings at the serving system and re-use them across requests from the same conversation. More concretely, one can save the KV-tokens in the KV cache belonging to a previous request and only process the new prompts for the next follow-up request.
Figure 3 demonstrates the heavy cost of prompt initiation phase under an artificial workload where each request generates 200 tokens and has varying conversation history sizes. As shown in the figure, the cost of recomputing the conversation history (solid blue line) causes the cost of the prefill phase to soon outgrow the generation phase.
3.2. Challenges
Limite GPU memory for caching.
LLM has very large model parameters, which results in large KV-tokens. For example, a 13 billion parameter GPT-3 model has 40 layers and a hidden size of 5120. Assuming the use of 16-bit half precision numbers, storing each KV-token takes 2 * 40 (layer) * 5120 (units/layer) * 2 (bytes/unit) = 0.78MB memory space. Under the current GPU memory limit, depending on the history lengths, only a few dozens or hundreds of conversation histories can be kept in the GPU. We must extend our cache space to use the more abundant CPU memory.
Cache management.
When using a multi-tier GPU-CPU cache, the serving system must be able to evict and restore cached data. The system has to make best-effort decisions to swap cached KV-tokens from GPU to CPU to make room for requests from other conversations. Later, it needs to recover the swapped out KV-tokens as fast as possible. When the system is under CPU memory pressure, some cached KV-tokens might need to be dropped. Therefore, we need a low overhead solution to manage memory space and swapping and deal with the case that some parts of the conversation history embeddings are no longer available in the system.
Handling non-contiguous KV cache
Figure 4. Multi-Query Attention with Non-contiguous Context Cache. Batch contains two requests. Left-hand side shows the Query representation for tokens being processed in each request. Right-hand side shows the context data (Key and Value representation) for each request in GPU.
In the multi-turn conversation setup, whenever a new requests comes in an existing conversation, the request’s new prompt, which in all likelihood has more than one input token, needs to be processed along cached KV-tokens as part of the prefill phase. For any cached KV-token that has been previously evicted to the CPU, there is no guarantee they are restored to the GPU in contiguous memory region next to existing KV-tokens cached in the GPU. Figure 4 provides an illustration of the computation pattern of attention. The batch contains request 0 which is currently performing generation using the last generated token, and request 1 which just arrives with a 7-token new prompt. To perform attention operation, request 0 needs to iterate over tokens in KV cache block 3 and 1, and request 1 needs to iterate over tokens in cache block 0, 4, 5, 2, 7. Existing high performance attention kernels, e.g. MemoryEfficientAttention (memeff, ), FlashAttention (flash, ), require inputs to reside in contiguous memory space and therefore cannot work with non-contiguous paged memory management. Although vLLM has developed the PagedAttention GPU kernel, it is designed to be solely used in the generation phase, and thus has the limitation that each request in the batch can have at most one input token. Hence, PagedAttention kernel cannot support prompt initiation for request 1 in Figure 4. And due to the design of single-query PagedAttention, no trivial way is available to extend it to support multi-query attention. An alternative approach might be to process the new prompt query in the same iterative manner as incremental decoding, one token a time, so that PagedAttention kernel can be applied. But this approach gives up the parallelization opportunity available in the prefill phase due to the extra query dimension, and will lead to significantly longer computation time for long input prompts.
4. System Design
In this section, we describe the system design of Pensieve. At a high level, we aim to save a conversation’s KV-tokens across multiple turns, if there is enough cache space to do so. To realize the potential performance benefits, we need to make cache swapping and missing token recomputation fast and efficient, by developing a few techniques to address the challenges in §3.
Figure 5. Overview of Pensieve
4.1. System Overview
Figure 5 shows the overall system architecture. Pensieve consists of a single scheduler and a worker engine to manage the GPU. The scheduler has two jobs: 1) it is responsible for batching requests for execution, and 2) it ensures that requests in the batch have sufficient GPU memory for execution. For 1), the scheduler performs fine-grained iteration-level batching (batchmaker, ; orca, ; vllm, ) so that a new request can join the batch with existing requests while the latter are in the process of performing autoregressive generation. For 2), the scheduler manages the allocation of slots in the KV cache and when to swap between the GPU and CPU KV cache.
The worker also has two jobs: 1) it invokes kernels to execute the LLM computation for a batch of requests. 2) it performs the actual data movements between the GPU and CPU KV cache based on a batch’s cache plan received from the scheduler.
4.2. A unified batch scheduler
Pensieve performs fine-grained iteration-level batching (batchmaker, ; orca, ; vllm, ). But our batching strategy differs from existing LLM serving systems. ORCA only batches non-attention operations and handles attention operations for each request individually (orca, ). By contrast, Pensieve batches both non-attention and attention operations, like vLLM (vllm, ). However, unlike vLLM which only forms a batch among requests in the same phase and thus processes each of the two types of batches in a separate model execution, Pensieve handles both the prefill phase and generation phase in a unified way. In other words, the Pensieve scheduler forms a batch of requests regardless of which phase they are in. In the same batch, some could be in their generation phase while others are in their prefill phase. Such unified batching is made possible by Pensieve’s multi-query attention kernel design (§4.4.1).
Pensieve’s iteration-level scheduler is “clocked” for action by the completion of a generation step. In particular, after each iteration of token generation, the worker returns any finished request that has either emitted the end-of-sentence token or reached maximum decoding length. The scheduler finds the next request in its wait queue to join the batch, if there is room (i.e. the number of total tokens in the batch is fewer than some pre-configured threshold). We use the simple first-come-first-serve scheduling policy when choosing which new requests to join the batch.
For unified batch formation, the scheduler concatenates the tokens to be processed from all requests. For each new request that has just joined the batch, it has a sequence of tokens corresponding to its prompt. For each existing request in the batch, it has a single token corresponding to the last generation step’s output token. By combining the prefill phase together with the generation phase, we avoid running separate small kernels and can thus improve GPU utilization.
4.3. KV cache management
Traditionally, the KV cache in GPU only serves as computation workspace. In Pensieve, the GPU KV cache also serves as storage space to cache the KV-tokens of recently completed requests of active conversations. Pensieve adopts a two-tier hierarchical caching strategy and uses the much larger CPU memory as the second tier cache space. The scheduler determines the caching plan and instructs the worker to perform the actual data movements between GPU and CPU.
The scheduler tracks the amount of free GPU KV cache slots left. Before handing off a batch of requests to the worker, the scheduler tries to ensure that any new request’s past KV-tokens will reside in the GPU and there is sufficient GPU memory for the execution. Specifically, if a request in the batch has any KV-tokens that have been swapped out to the CPU memory or dropped, the scheduler determines how much additional GPU KV cache slots are needed to swap in or re-calculate those absent KV-tokens. If there is sufficient space, the scheduler instructs the worker to perform the necessary allocation and to start swapping those KV-tokens from the CPU as part of the batch’s cache plan.
Pipelined KV cache recovery.
The scheduler does not wait for a request’s KV-tokens to be fully swapped in from the CPU before handing it off to the worker for execution. Rather, we follow the pipelined approach (pipeswitch, ) to overlap computation with data transfer. Specifically, we exploit the fact that an LLM model has many layers and each layer’s KV-token is only used in this layer’s self attention calculation. Instead of waiting for all layers to finish data transfer before starting the execution, we initiate the transfer layer by layer and start model computation at the same time. The worker uses GPU events to preserve data dependency: it only starts a layer’s self-attention kernel once that layer’s KV-tokens have been fully copied to the GPU. Pipelining transfer with computation allows us to hide the swap-in latency.
Ahead-of-the-time swapping and the eviction policy.
Since Pensieve tries to preserve KV-tokens in the GPU for reuse across multi-turn conversations, the scheduler does not immediately release a request’s GPU cache slots as soon as it finishes, unlike existing systems (vllm, ). Instead of waiting until the GPU cache has run out, the scheduler asks the worker to copy (aka swap out) selected KV-tokens to the CPU if less than a threshold (e.g. 50%) of the GPU cache slots are available. The corresponding GPU memory is not immediately released until the scheduler later decides to allocate the same slots to another conversation.
The scheduler selects which conversation to swap out by balancing two factors: 1) the number of KV-tokens the conversation has in the GPU cache. The larger the size is, the more free space can be released. 2) the amount of time since the conversation has last sent a request, as also used in LRU. The scheduler weighs both factors by multiplying them to get a score for each conversation and preferentially swaps out those with higher scores. Our swapping is fine-grained; the scheduler only evicts half of a conversation’s KV-tokens at a time, specifically the trailing half. The scheduler iteratively decides which conversation to swap out based on their scores until the GPU cache is no longer under memory pressure. The scheduler uses the same eviction policy for the CPU cache. Tokens evicted from the CPU are simply dropped.
Handling dropped tokens.
If a request has some of its KV-tokens dropped due to CPU memory pressure, we resort to recomputation to handle such dropped tokens. Figure 6 illustrates the most general scenario for the whereabouts of a new request’s KV-tokens. As we always evict cached tokens from the trailing end of a conversation, the GPU cache generally holds the request’s early tokens, and the CPU cache holds the middle, while the rest may have be dropped. The scheduler swaps in those tokens cached in the CPU. For dropped tokens, the scheduler will fetch their corresponding raw text tokens from the conversation history saved in a persistent store (Figure 5). These retrieved raw tokens will be merged into (i.e. prepended to) the new request’s prompt and become part of the batch’s input tokens, as shown in Figure 6. Thus, the dropped tokens will be recomputed as part of the prefill phase for the new request. The policy of evicting/dropping from a context’s trailing end makes such merging possible.
Figure 6. Recompute KV-tokens dropped by the CPU cache as part of the new request’s prefill phase.
Suspending requests during generation
In spite of ahead-of-time eviction, the scheduler might still encounter scenarios when the generation phase runs out of GPU cache since requests’ decoding lengths are not known a priori. In this situation, the scheduler suspends some requests’ execution by taking them out of the current batch, swaps out their corresponding KV-tokens to the CPU and puts them back in the waiting queue. It chooses which request to suspend according to the descending order of their arrival time. As suspension causes increased latency (due to waiting for the swap-out), we try to avoid it by conservatively reserving 10% of GPU cache slots for the execution of existing requests that are in the generation phase. In other words, the scheduler stops adding new request into the running batch unless there are more than 10% free GPU cache slots.
Figure 7. Single-query attention vs multi-query attention, softmax and rescaling omiited. $d$ is dimension of the embedding vector. red shaded area in multi-query attention score represents causal masking. $K$ and $V$ are logically contiguous, but physically placed at random locations when paged memory management is used.
4.4. Multi-query attention for non-contiguous cache
How to combine existing KV-tokens in GPU cache with those just swapped in from the CPU? The most naive solution is to allocate a contiguous memory region in the GPU to hold them both. However, doing so would incur expensive memory copying, since KV-tokens are large. A more promising solution is to allocate separate space only for those swapped-in KV-tokens and to design an attention kernel implementation that can handle non-contiguous memory in its KV cache.
vLLM’s PagedAttention kernel can handle non-contiguous KV cache in the generation phase (vllm, ). However, for the prefill phase, it still uses existing kernels which require all KV-tokens to reside in contiguous memory. We cannot use PagedAttention because it assumes each request in a batch has exactly one input token, which is the case for generation but not prefill. Thus, we refer to PagedAttention as a single-query attention kernel because it computes the attention scores between a single input token’s query representation (Q) and the KV cache. We need to build a multi-query attention kernel which supports performing attention between the the query representations (Q) of multiple input tokens per request and the KV cache over non-contiguous memory.
Our new kernel is similar to that of vLLM to the extent that both kernels need to support loading KV cache from non-contiguous GPU global memory into on-chip shared memory. Their main difference is illustrated in Figure 7. As vLLM’s kernel performs attention between a request’s single input token and the existing KV-tokens, its underlying computation can be described as two matrix-vector multiplication operations, as shown in Figure 7  [^1] The batched version of PagedAttention performs batched matrix-vector multiplications.. In contrast, our kernel handles multiple input tokens for each request, computing attention scores between all pairs of input tokens and the conversation’s existing KV-tokens. Therefore, its underlying computation can be described as two matrix-matrix multiplication operations, and the batched version of our kernel performs batched matrix-matrix multiplications. Like vLLM, we fuse the two multiplication operations according to (flash, ). Because of the additional dimension in the Q tensor, our kernel has more parallelization and tiling opportunities on the GPU. However, care must be taken to handle the new challenge that different request in the batch have different number of input tokens.
When multiple input tokens of a request are handled together in a kernel, we need to apply causal masking so that a later input token does not “attend” to the earlier tokens. This requires setting a corresponding upper triangular part of the attention score matrix to 0, as shown by the red shaded area in Figure 7. Causal masking is not needed for the single-query attention kernel since it only processes one input token. We fuse the causal masking operation inside the multi-query attention kernel to avoid materializing the intermediate attention score matrix (flash, ).
Matrix-matrix multiplication kernels are much more complex than those for matrix-vector multiplication because they typically use sophisticated algorithms to extract additional data reuse opportunities with GPU’s on-chip shared memory and to leverage GPU’s tensor core primitives. Therefore, instead of trying to extend vLLM’s PagedAttention kernel for multi-query attention, we base our implementation on an existing multi-query kernel from PyTorch and extend it to handle non-contiguous KV cache. Our kernel uses the high performance thread-block level matrix-matrix multiplication provided by NVIDIA Cutlass template library (cutlass, ). Unlike PagedAttention, our kernel does not group multiple KV-tokens into blocks. Instead, each KV-token can potentially reside in a different memory location. Our kernel accepts a list of slot indices to gather KV-tokens from scattered memory locations. Such flexibility helps reduce internal fragmentation compared to the paged design.
4.4.1. Unifying the prefill and generation phase
As discussed in §4.2, in Pensieve’s batch formation, new requests in their prefill phase can be grouped together with existing requests in their generation phase. This is enabled by our multi-query attention kernel, because single-query attention performed in the generation phase can be treated as a special case of the multi-query attention with query size equal to 1.
More concretely, Pensieve’s scheduler concatenates the input tokens to be processed from all requests, regardless of whether they correspond to prompts of new requests or to last step’s output token from existing requests. Some auxiliary data structure is maintained to keep track of each request’s corresponding region. During execution by the worker, these batched input tokens are fed through linear layers to generate each token’s QKV representations. Newly generated KV-tokens are stored into the allocated slots in the GPU’s KV cache. Then the worker applies our multi-query attention kernel to produce the output tokens for all requests.
5. Implementation
We have implemented our prototype serving system Pensieve with $\sim$5K lines of C++/CUDA code. Pensieve relies on PyTorch C++ API front-end to perform GPU Tensor memory management and to execute Large Language Models. Based on PyTorch’s fused memory efficient attention, we develop our own multi-query fused attention kernel using NVIDIA Cutlass library to support performing attention with KV-tokens that reside in non-contiguous memory.
5.1. Optimization: Prioritize Data Transfer to GPU over eviction to CPU
Although PCIe sends data full-duplex bidirectionally, in practice, we found that when CPU-to-GPU transfer is overlapped with GPU-to-CPU transfer, there is a significant throughput drop (18-20%) in both directions. Similar issues have been reported [^2] https://forums.developer.nvidia.com/t/data-transfers-are-slower-when-overlapped-than-when-running-sequentially/187542. Since Pensieve performs KV-token backup ahead of time, there is no urgency to finish the transfer right away. In order to prevent evicting from slowing down swapping in past KV-token, we set up a blocking mechanism. Whenever worker plans to issue GPU-to-CPU copy, it first checks if there is any swap-in task in the system, and waits until swap-in task is done. Although this conservative approach does not fully utilize duplexed PCIe bandwidth, due to the existence of generation phase, we find that this optimization performs well and we never run into the situation that the backup can’t catch up.
6. Evaluation
In this section, we show the evaluation results of Pensieve.
Model size 6.7B 13B
# layer 32 40
# hidden 4096 5120
# head 32 40
Layer MFLOPs Linear 402.65 629.15
Attention 33.55 41.94
Total 426.20 671.09
Table 1. Evaluated GPT-3 model configurations, and approximate amount of computation (FLOPs) in each Transformer layer normalized by the number of input sequence tokens.
ShareGPT LMSYS
# conversations 15553 79166
Average # turns 7.58 7.73
Average input length 31.26 22.49
Average output length 180.44 100.67
Table 2. Dataset staticstics
(a) GPT3-6.7B, ShareGPT
(b) GPT3-13B, ShareGPT
(c) GPT3-6.7B, LMSYS
(d) GPT3-13B, LMSYS
Figure 8. 90 percentile normalized latency vs throughput graph for GPT-3 model on ShareGPT and LMSYS datasets.
6.1. Experimental Setup
System Environment
We evaluate Pensieve on a Microsoft Azure NC24ads_A100_v4 instance, which is equipped with 24 non-multithreaded AMD EPYC Milan processor cores, 220 GB system memory, and one NVIDIA A100 PCIe GPU with 80GB memory. We configure each system to allocate 40GB GPU memory to KV cache for fair comparison. Pensieve and all baselines use CUDA version 11.8 and PyTorch 2.0.
Model
We use the popular GPT-3 (gpt3, ) as our benchmark model and evaluate two configurations listed in the GPT-3 paper: 6.7 billion parameters and 13 billion parameters. Detailed model hyper-parameters can be found in Table 1. For all experiments, 16-bit half-precision float format is used for both model parameters and intermediate hidden representations. Model parameters are randomly initialized.
Dataset
We evaluate Pensieve using two datasets collected from AI-powered ChatBot: ShareGPT (sharegpt, ), a dataset containing user-shared ChatGPT conversations [^3] We used this version: https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, which performed additional filtering like removing HTML tags and non-English conversation., and LMSYS-Chat-1M (lmsys, ). Since we target multi-turn conversation scenario, we only keep conversations with at least 5 turns. We also limit the maximum context size to 2048 (larger context size studied in §6.3). Statistics about datasets are listed in Table 2. We extract the input and output lengths of each conversation turn from these datasets to construct requests. To ensure the same computation amount, we ask each system to ignore any early stop criteria, e.g. end-of-sentence (EOS) token, and keep generating until the ground-truth output length is reached.
Workload
Since the dataset does not provide timestamps for when each user prompt is received, we sample arrival time for each request from Poisson distribution under different request rates. We maintain the causal dependency for requests belonging to the same conversation: a new user prompt is only sent to the system after earlier ones from the conversation have already received their response. In addition, in real-life chatting scenarios, it takes time for users to reply. We assume each generated output token adds 100 milliseconds to such user response time. For example, if 200 tokens are generated for a conversation, then within the next 20 seconds, no new request will come from this conversation. Such response time is set to have a minimum value of 5 seconds.
Baseline
We evaluate our system against the state-of-the-art vLLM serving system (vllm, ) (version v0.2.0), which uses a stateless serving API. For each request, the user needs to concatenate chat history with the new user question as input prompt. Since vLLM is a PyTorch-based execution engine with PagedAttention kernel integrated, and has no perpetually serving loop to drive the engine to continue execution, we implement such a driver that adds newly arrived requests into vLLM job queue and invokes the engine execution until all requests are fully processed.
We also implement a variant of Pensieve called Pensieve (GPU cache) that does not use CPU memory and only saves past KV-tokens in spare GPU memory spaces to reuse them in the future.
Performance Metric
Pensieve is designed to optimize both peak throughput and latency of serving LLM in conversational scenarios. Following prior work (orca, ; vllm, ), we measure the 90 percentile normalized latency and the achieved request serving throughput. Normalized latency is calculated as the end-to-end request serving latency divided by the number of output tokens.
6.2. End-to-end Pensieve Performance
Figure 8 is the 90 percentile latency (normalized by request output length) vs throughput graph of the end-to-end serving performance of Pensieve and baselines. When compared against vLLM, for GPT3-6.7B and GPT3-13B, Pensieve is able to achieve a throughput gain of 1.51x and 1.65x on ShareGPT dataset, and 1.84x and 1.95x on LMSYS dataset. At the point when vLLM baseline is close to its peak throughput, Pensieve is able to reduce 90 percentile normalized latency by about 60-75%.
The throughput gain on LMSYS dataset is in general larger than ShareGPT dataset. The reason is that requests in LMSYS dataset have on average smaller output length (see Table 2), and as a result, the amount of time spent on generation phase is shorter. Since Pensieve focuses on optimizing prompt prefill phase by avoiding recomputing past KV-tokens, a dataset with a relatively longer prefill phase favors our solution.
Experiments using GPT3-13B achieve about 12% higher throughput gain on both datasets when compared against GPT-6.7B. This is due to the fact that as the model dimension increases, the computation amount increases faster than the context cache size that we need to transfer from CPU. The bottom part in Table 1 lists the computation amount by major operations in each Transformer Layer. Linear layer operations  [^4] Despite having more FLOPs in absolute number, linear layer operators can make use of highly efficient NVIDIA Tensor Cores. grows quadratically with respect to the growth of hidden size. When the model hidden size is increased from 4096 to 5120, which amounts to 25% more data to copy, linear operation’s computation amount increases by around 50%. Therefore, recomputing full context from scratch becomes more costly. With the trend of increasingly larger models, we expect to see Pensieve becoming even more beneficial.
Pensieve (GPU cache), the yellow line in Figure 8, demonstrates the performance of using GPU cache only without evicting to CPU. This approach utilizes unoccupied GPU space to keep past KV-tokens of conversations for as long as possible, and only discard them when GPU KV cache is out of memory. When the system is serving under a low request rate, GPU-side caching only may also deliver a similar serving latency of Pensieve. However, as workload rates go up, more conversations are being served at the same time, and GPU memory space exhausts quickly. Eventually, at a higher rate, this approach behaves similarly to recomputing past KV-tokens from scratch.
6.3. Larger maximum context size
(a) ShareGPT
(b) LMSYS
Figure 9. 90 percentile normalized latency vs throughput graph for GPT3-13B model with maximun context size 16384
In scenarios where the conversation lasts lots of turns, or the user feeds a long article for the AI agent to read and answer questions, context size may exceed the commonly used 2048 limit. Figure 9 shows how longer context affects the performance. From the figure, with maximum context size 16384, the throughput gain Pensieve can achieve over the baselines decreases. On one hand, longer conversations need more memory to store. On the other hand, longer conversations tend to have more turns, last longer, and have to be stored longer, making it more likely for the system to experience KV-token drop and recomputation.
6.4. Efficiency of multi-query attention kernel
Figure 10. Execution time of Pensieve’s multi-query attention kernel over non-contiguous context memory with different context size. Measured with batch size 32 and query size 8.
Figure 10 shows the advantage of Pensieve’s multi-query attention kernel when used to prefill KV-tokens of prompts from new requests with their past KV-tokens. In this microbenchmark, a batch of 32 requests perform attention between a prompt of 8 query tokens and different numbers of past KV-tokens stored in non-contiguous GPU memory. As described in §3.1, existing attention kernels are not directly applicable. We compare against two straw-man implementations: 1) allocate additional contiguous GPU memory space to copy out past KV-tokens and then apply a fused attention kernel (orange bar), and 2) apply multiple rounds of vLLM’s single-query PagedAttention to process one token from the prompt at a time (green bar). We also present the performance of the ideal situation which assumes that past KV-tokens are already in contiguous memory space (blue bar). From Figure 10, both straw-man solutions add significant overhead compared to the ideal performance. Copy out requires extra memory space and gathers past KV-tokens from non-contiguous space. Such cost is proportional to the number of past KV-tokens. Applying multiple rounds of PagedAttention, on the other hand, gives up parallelization opportunities on prompt tokens, leading to linear execution time with respect to the number of tokens in the prompt. Pensieve’s kernel shows slightly better performance than the ideal baseline because we offload auxiliary data computing (like calculating the cumulative sum for the sequence length of a batch) to CPU. Since each transformer layer in the model shares the same caching plan, these auxiliary data can be reused by all layers.
7. Discussion
Distributed Execution
For large LLM model that needs to be sharded across multiple GPUs, Pensieve follows Model Parallelism (also called Tensor Parallelism) pattern of Transformer layers proposed by NVIDIA Megatron-LM (megatron, ). With model parallelism, each GPU worker executes exactly the same computation and therefore, the same context cache arrangement can be shared by all workers.
Distributed version Pensieve is not fully vetted yet. We will provide concrete evaluation results for multi-GPU and larger models in next version of this paper.
8. Related Works
8.1. Systems Tailored for Serving Language Models
Many systems have been designed to serve language models with high performance. For example, systems like FasterTransformer (fastertransformer, ), LightSeq (lightseq, ), DeepSpeed (deepspeedinference, ), etc., adopt incremental decoding to avoid duplicated computation, provide better fused kernels and quantization support, and improve the efficiency of decoding algorithms like beam search and various sampling algorithms.
A key challenge presented by language model serving is that requests may have different input prompt sizes, and the number of output tokens to generate is unknown beforehand until stopping criteria is met. Most systems handle this variable length issue by padding each request in the batch to the same lengths. In order to make more efficient batching decisions, BatchMaker (batchmaker, ) proposed to use finer batching granularity and perform scheduling at the granularity of the token. ORCA (orca, ) further extended this idea to Transformer-based models by adding support to attention operations.
Among all systems serving language models, vLLM (vllm, ) is the most closely related one to this work, as it focuses on LLM KV cache memory management. vLLM addresses memory fragmentation caused by variable request prompt size and unknown generation size by adapting virtual memory paging solution from the operating system. It divides memory into pages, each holding a fixed number of tokens. Instead of allocating memory to maximum context size, vLLM allocates physical GPU memory on demand and enables memory sharing among requests.
8.2. Techniques Targeting Limited GPU Memory
Many techniques address GPU’s limited memory space, including swapping, recomputation, unified memory, etc.
Swapping / offloading
SwapAdvisor (swapadvisor, ) performs GPU-CPU swapping for DNN training. It analyzes dataflow graphs and decides operator execution order, memory allocation, and the swapping plan. Zero-Offload (zerooffload, ) offloads optimizer states and gradients to CPU for training LLM with a single GPU. DeepSpeed-ZeroInference (deepspeedinference, ) and FlexGen (flexgen, ) further adapt the offloading techniques to serve LLM with a weak GPU. DeepSpeed-ZeroInference offloads entire model weights to CPU or NVMe memory. FlexGen offloads currently unused model weights, activation or KV cache to CPU memory or disk. But as a result of frequent data movement and high disk latency, these two systems require a large batch size so that computation amount is large enough to hide the offloading latency. Therefore, they mainly target latency-insensitive applications.
Recomputation
In training, model activation recomputation (recompute, ) is a commonly used technique to reduce memory footprint. A subset of activations with relatively low computation cost is selected to be dropped after being used in model forward pass and recomputed on demand in backward pass. Checkmate (checkmate, ) further formalizes the trade-off between time and memory usage as tensor rematerialization optimization problem and finds the optimal schedule using MILP solver. (reducing, ) exploits the Transformer layer structure to only selectively recompute attention module which has larger memory consumption but low number of floating-point operations to improve the cost efficiency of trading recomputation time for memory space. But in inference setting, we only preserve the Key Value embedding of tokens for incremental decoding and all other activations are discarded. KV embeddings are very costly to recompute, and therefore, we opt to swap them to CPU.
Unified Memory and Direct Host Access
Moving LLM KV cache between CPU and GPU memory can also be achieved in a transparent manner through Unified Memory (unifiedmemory, ), which, similar to CPU virtual memory, allows users to oversubscribe GPU memory and automatically triggers memory page migration between CPU and GPU using a page fault mechanism. However, CUDA Unified Memory may cause longer latency due to the overhead of page migration. DeepUM (deepum, ) addresses this issue by developing better prefetching techniques. HUVM (huvm, ) further provides unified multi-GPU virtual memory and allows transparently storing data on neighboring GPUs to utilize the high-bandwidth interconnect among GPUs to reduce access latency caused by host memory access via slower PCIe.
Another line of works (direct, ; deepplan, ) utilizes Direct Host Access feature of modern GPU, which allows GPU kernel to directly read from host memory. The main difference compared to CPU to GPU memory copy (explicit CUDAMemcpy or implicit Unified Memory) is that the data transfer to GPU is immediately consumed by the GPU thread requesting it. This brings better overlapping of data transfer and computation because part of threads in GPU kernel can start to execute earlier without waiting for the entire transfer to finish as long as the data these threads requested are ready in GPU.
Neither Unified Memory nor Direct Host Access is the most efficient choice to solve our KV cache management problem, because their memory transfer is only triggered when data is accessed by a GPU kernel, which unnecessarily gives up the opportunity to fetch ahead of time. Hence, we decided to manage data movement explicitly. Furthermore, for Direct Host Access, since the copied data is not explicitly stored in GPU memory, it has to be loaded from CPU again if it is repeatedly used. In our workload, the context cache is not only used to prefill the prompt tokens but also in every generation step, which makes Direct Host Access significantly less efficient.
8.3. Techniques that speed up LLM inference
Kernel Fusion Deep Neural Networks often have a lot of memory-bound operations, e.g. element-wise operators, that read a large amount of data while performing little computation. A common technique to improve GPU efficiency is kernel fusion (xla, ; tvm, ), which fuses multiple kernels as one single kernel to avoid storing intermediate data into GPU global memory and later loading them back. Attention in LLM can also benefit from kernel fusion. Memory efficient attention (memeff, ) and Flash Attention (flash, ) fuse all attention-related operations into one kernel to avoid materializing the attention score tensor. The attention kernel we implemented in Pensieve is a generalization of above mentioned fused kernels, with support to efficiently attend new tokens with past KV-tokens stored in non-contiguous memory. Recently, Flash-decoding (flashdecode, ) further parallelizes the processing of KV-tokens for single query token attention, which can significantly speed up the generation phase. With prefill phase taking up a larger portion of total computation, Pensieve’s optimization for prefill phase may become even more beneficial.
Quantization Quantization techniques use lower precision floating point or integer representation to speed up execution and reduce memory footprint (nagel2021white, ; gholami2022survey, ). 16-bit half precision is widely adopted in LLM inference and mixed precision training (mixed, ). And recently, several works (dettmers2023case, ; qlora, ; flexgen, ) have shown that quantizing to 4-bit can still maintain good LLM inference quality. In this work, we focus only on 16-bit half-precision to deliver unchanged inference results. But in general, quantization-based methods are also applicable to our system.
Speculative Decoding Recently, speculative decoding (speculative:chen, ; speculative:leviathan, ) is proposed to speed up the generation phase in LLM inference. Instead of running a costly LLM model to sequentially produce 1 output token in each model iteration, speculative techniques use a smaller and faster model to generate a draft of a series of tokens, which get evaluated by a larger and more powerful LLM model later with more efficient parallel execution. The adoption of speculative decoding significantly reduces the execution time of generation phase, which makes Pensieve’s optimization of prefill phase in multi-turn conversation serving more important.
8.4. Other General DNN Serving Systems
There are many general-purpose systems, like TensorFlow Serving (tfserving, ), Clipper (clipper, ), NVIDIA Triton Inference Server (triton, ), Nexus (nexus, ), InferLine (inferline, ) that serve as scheduling component of the serving system. While being mostly model-agnostic and execution backend agnostic, these systems apply general system techniques like batching, caching, and software pipelining to serve Deep Learning models, and also properly provision compute resources to improve overall cluster efficiency. On the other end of the spectrum, several works target model-less serving. They aim to provide inference as a service and automatically select models to meet accuracy and latency requirements. For example, INFaaS (infaas, ) and Tabi (tabi, ) serve requests with small model and only re-route to larger models when the output confidence score is low.
9. Conclusion
In this work, we identify that a major inefficiency in serving Large Language Models for multi-turn conversations is the recomputation of history context. We introduce Pensieve, a stateful LLM serving system that preserves context data in a multi-tier GPU-CPU cache and we develop a new GPU attention kernel that supports performing attention between requests’ new multi-token input and their saved context stored in non-contiguous GPU memory. Experiments show that Pensieve improves serving throughput 1.51-1.95x and reduces latency by 60-75% compared to the state-of-the-art system vLLM.